OpenStudy (18jonea):

Use elimination to solve the system of equations given by –7x + 3y = 14 and 7x – 2y = -11. (−5/7,3) (−1 5/7,−35) no solution infinitely many solutions

Vocaloid (vocaloid):

start by adding the two equations together

OpenStudy (18jonea):

0 + y =3 @Vocaloid

OpenStudy (18jonea):

@Cuanchi

Vocaloid (vocaloid):

good! that means y = 3 now, substitute y = 3 back into one of the equations to find x

OpenStudy (18jonea):

3 and -5/7 @Vocaloid

Vocaloid (vocaloid):

good!

OpenStudy (18jonea):

Which of the following describes the system of equations 2x - y + 4 = 0 and 4x - 2y + 7 = 0? would this be inconsistent

OpenStudy (18jonea):

@Vocaloid

Vocaloid (vocaloid):

hm, I don't think so

Vocaloid (vocaloid):

actually, never mind, you're right

Vocaloid (vocaloid):

it's inconsistent since the slopes are the same, and the y-intercepts are not

OpenStudy (18jonea):

ok thank you